OpenStudy (anonymous):

Solve the system by substitution -x-y-z=-8 -4x+4y+5z=7 2x+2z=4

OpenStudy (adamaero):

"substitute" one equation if for the other (you have to solve for one variable first: this can be done with either the first or last equation)

OpenStudy (anonymous):

so can you show me how to do that

OpenStudy (adamaero):

-x-y-z = -8 -x = -8 + y +z x = 8 - y - z plug (8 - y - z) for x

OpenStudy (adamaero):

actually you have to do this kind of thing twice